OpenStudy (anonymous):

Find the point-slope form of the line that passes through given the point and has the given slope.Include your work in your final answer. Type your answer in the box provided or use the upload option to submit your solution. (0, 0) m = 2

OpenStudy (dayakar):

point (x1,y1) =(0,0) slope (m) =2 given y-y1=m(x-x1) y-0=2(x-0) y=2x

OpenStudy (anonymous):

thank you

OpenStudy (dayakar):

it's allright